Two equal deposits made 20 and 21 years ago, respectively, will allow a retiree to withdraw

$10,000 now and $10,000 per year for 14 more years. If the account earned interest at 10% per

year, how large was each deposit?
Answer #1

Present value of both withdrawals ($) = 10,000 + 10,000 x P/A(10%, 14) = 10,000 + 10,000 x 7.3667 = 10,000 + 73,667

= 83,667

Let each deposit be $D. Then for equivalence,

D x F/P(10%, 20) + D x F/P(10%, 21) = 83,667

D x 6.7275 + D x 7.4002 = 83,667

D x (6.7275 + 7.4002) = 83,667

D x 14.1277 = 83,667

D = $5,922.20